I would really appreciate some help with this one.
The half life of a radio active substance is 125 days. How many days will it take for 82% of the substance to decay?

Answer by scott8148(6628)

At = Ao (1/2)^(t/h) ___ At is amount at time t ; Ao is initial amount ; t is the time ; h is the half life

At / Ao = (1/2)^(t/h)

1 - .82 = (1/2)^(t/125)

log(.18) = (t/125)log(1/2)

[log(.18)] / [log(1/2)] = t / 125